Dustin is considering an investment that will pay $3000 a year for 5 years, starting 1 year from today (which is normal). How much should Dustin pay for this investment today if he wishes to earn a 7 percent annual rate of return?
Answer #1

Present value of Annuity = A*[(1-(1+r)-n)/r]

Where

A - Annuity payment = 3000

r - rate per period = 7%

n - no. of periods = 5

Present value of Annuity = 3000* [(1-1.07^-5)/.07]

= 3000* [(1-0.71298617948)/.07]

= 3000*0.28701382052/.07

= 3000*4.100197436

= $12300.59